OpenStudy (anonymous):

solve this system of equation using subtraction x=4y 3x-2y=20

OpenStudy (anonymous):

Substitute 4y for x in the 2nd equation. 3(4y)-2y = 20 12y-2y = 20 10y = 20 y = 2 Now, substitute 2 for y in the 1st equation. x = 4(2) x = 8 So, the solution is (8,2)

OpenStudy (anonymous):

what about x + 3y = 2 2x - 3y = 22